# EgoDTM: Towards 3D-Aware Egocentric Video-Language Pretraining

### v1

@@ 4.4 More Analysis
Method mAP @(0.5:0.95) map @0.50 hand-AP @(0.5:0.95) object-AP @(0.5:0.95)
Upperbound [12] 60.70 73.31 90.87 30.53
Our Pipeline 43.02 54.88 61.01 25.03
Table 6: Evaluation of our HOI mask generation pipeline on HOI segmentation benchmark VISOR [12] in kitchen environment.
@@
Quantitative and qualitative analyses of the HOI mask generation pipeline. To evaluate the reliability of the HOI detector and SAM2 model in the egocentric domain, particularly for hand-object interactions, we evaluate the image-based HOI segmentation on the VISOR dataset [12]. The results are presented in Table 6. The supervised training model on VISOR serves as the upper bound for performance comparison. While our generation pipeline does not fully reach the upper bound, it demonstrates high-quality segmentation results. Notably, hand segmentation is significantly lower than the upper bound, likely due to inaccuracies in HOI detection prompts and challenges in segmenting hand-object interactions with SAM2 in cluttered egocentric backgrounds.  Figure 5 presents that our data construction pipeline is useful even when HOI regions are small while the background is noisy on Ego4D.
Visualization of predicted depth. In Figure 3, we visualize predicted depth maps in unseen environments. Despite operating with low-resolution inputs, EgoDTM effectively estimates depth, yielding results close to those of DAv2. Notably, there is a clear correlation between the depth maps and the hand-object interaction regions.
@@
Figure 5: Example of generalizable data construction. For better visualizations, we blur the background to highlight HOI regions.
5 Conclusion
In this work, we present EgoDTM, a novel egocentric 3D-aware video-language model. EgoDTM integrates dual transformer encoders with a lightweight 3D-aware depth decoder, trained using video-text contrastive learning and depth estimation objectives. To enable large-scale pretraining, we generate millions of depth maps and spatially enriched captions by leveraging foundation models. The captions are enhanced through a detect-track-generate pipeline specifically tailored for egocentric videos. EgoDTM demonstrates intuitive generalization in estimating depths in unseen environments. Extensive experiments across diverse benchmarks—spanning short video understanding, 3D understanding, and long video understanding—validate the effectiveness of our approach.
Limitations. While EgoDTM achieves strong performance in egocentric scenarios, its impact on broader domains remains limited. We will explore pretraining 3D-aware video representations for general domains in the future.

### v2

@@ 4.4 More Analysis
[…]
Quantitative analyses of the HOI mask generation pipeline. To evaluate the reliability of the HOI detector and SAM2 model in the egocentric domain, particularly for hand-object interactions, we evaluate the image-based HOI segmentation on the VISOR dataset darkhalil2022visor. The results are presented in Table 12. The supervised training model on VISOR serves as the upper bound for performance comparison. While our generation pipeline does not fully reach the upper bound, it demonstrates high-quality segmentation results. Notably, hand segmentation is significantly lower than the upper bound, likely due to inaccuracies in HOI detection prompts and challenges in segmenting hand-object interactions with SAM2 in cluttered egocentric backgrounds.
Qualitative analyses of the HOI mask generation pipeline. We firstly compare inconsistent HOI bounding boxes from HOID hoidetector (frame-by-frame detection) with spatial-temporal HOI masks generated by combining HOID hoidetector and SAM2 sam2 in Figure 9. The HOI boxes detected by HOI detector often lose tracks, since they are detected by an image-based model. Our pipeline achieves consistent HOI tracking across frames, offering more precise HOI labels.  Figure 9 presents that our data construction pipeline is useful even when HOI regions are small and the background is noisy.
5 Conclusion
In this work, we present EgoDTM, a novel egocentric 3D-aware video-language model. EgoDTM integrates dual transformer encoders with a lightweight 3D-aware depth decoder, trained using video-text contrastive learning and depth estimation objectives. To enable large-scale pretraining, we generate millions of depth maps and spatially enriched captions by leveraging foundation models. The captions are enhanced through a detect-track-generate pipeline specifically tailored for egocentric videos. EgoDTM demonstrates intuitive generalization in estimating depths in unseen environments. Extensive experiments across diverse benchmarks, spanning short video understanding, 3D understanding, and long video understanding, validate the effectiveness of our approach.
Discussions. While EgoDTM demonstrates strong performance in egocentric hand-object interaction scenarios, its generalization to broader indoor scenarios remains limited. Further exploration may include integrating 3D-aware visual encoders into multimodal large language models to enhance spatial awareness. Moreover, pretraining large-scale spatial-aware egocentric models with richer 3D signals, as explored in VGGT wang2025vggt, remains a promising yet challenging direction.
Acknowledgments
This work was partially supported by the Beijing Natural Science Foundation (No. L233008) and the Outstanding Innovative Talents Cultivation Funded Programs 2024 of Renmin University of China.
We provide additional details, extended experimental results, and further discussion in this supplementary material, including:
  • Implementation details: background of foundation models (Section A.1), dataset details (Section A.2), and experimental details (Section A.3).
  • Additional quantitative results: analyses and evaluations (Appendix B).
  • Additional qualitative results: Examples and visualizations to complement the main results (Appendix C).
  • Discussion on related works: Insights and comparisons with prior research (Appendix D).
Appendix A Implementation Details
A.1 Background of Foundation Models
HOI Detector hoidetector. HOID is a robust system for detecting human hands and their interacting objects in images. It is built upon Faster-RCNN ren2016fasterRCNN, pretrained on 100K image dataset with hand-object interaction annotations, including hand bounding box, interacting object bounding box, hand side (left or right), hand contact state (e.g., no contact, self-contact, other person contact, contact with portable object, or contact with a non-portable object). To enhance its capabilities, HOID is further trained with an additional 42K egocentric data samples, enabling improved understanding of HOI from egocentric view. We leverage HOID to generate spatial HOI boxes, hand side, hand contact state for each video, from 12 uniformly sampled frames per video. However, HOID often produces temporally inconsistent box predictions across adjacent frames. To address this, we apply a robust image and video segmentation model to refine the detection results, ensuring greater consistency and accuracy.
Segment Anything 2 sam2. SAM2 is a versatile segmentation model capable of segmenting objects in both images and videos according to a given prompt, such as a point, box or mask, with remarkable efficiency. It is trained on a large-scale SA-V dataset, comprising 50.9K videos and 35.5M high-quality masks. SAM2 employs a hierarchical image encoder and a memory mechanism to handle streaming frame input. In our approach, SAM2 is utilized to generate the spatial-temporal consistent HOI masks by leveraging prompts derived from HOID outputs.
Depth Anything 2 depthanythingv2. DAv2 excels in monocular depth estimation, offering fine-grained details, strong generalization and efficient inference. It is built upon the pretrained visual foundation model DINOv2 oquab2023dinov2 and a depth decoder DPT dpt. Pretrained on 595K synthetic images and 62M pseudo-labeled real images, DAv2 exhibits strong out-of-domain generalization. In our work, we use DAv2 to generate depth maps for eight frames sampled from egocentric videos, serving as supervision signals. Following the recommendations of DAv2, we employ the DAv2-Large variant, which produces more spatial-temporal consistent depth maps.
DeepSeek-LLM deepseekllm. We employ the LLM to interpret HOI information and enrich textual descriptions with shape and movement details. DeepSeek-LLM demonstrates exceptional ability to follow instructions and comprehend HOI mask prompts. Specifically, we use the DeepSeek API to access their 200B-parameter LLM to facilitate these tasks.
A.2 Dataset Details
Ego4D grauman2022ego4d. Ego4D contains 3,670 hours of egocentric videos with dense narrations, covering diverse scenarios and activities from worldwide. Each narration is timestamped and paired with a free-form sentence. Following the approach in  lavila, we construct 4M video-text clip pairs for pre-training, with an average clip length of 1 second (±0.9). In our text enrichment process, we only keep those hand-object interaction clips performed by the camera wearer, where the text begins with ‘#C‘ (denoting the wearer) and then follows HOI-related verbs and nouns. This strategy excludes clips that record other people’s activities, such as multi-person interactions (Ryan2023EgocentricAA) where the text begins with ‘#O‘, and the videos like ‘#C C walks away‘. For the natural language query task, it comprises 1,659 untrimmed videos, each averaging 500 seconds in duration. On average, each video contains 12 clip-query pairs. Following the official split from grauman2022ego4d, we use 11,291 queries for training and 3,874 for validation.
Epic-Kitchens epic-kitchens-100. Epic-Kitchens-100 (EK-100) consists of 100 hours of egocentric cooking videos divided into training (67,217 clips), validation (9,668 clips), and testing (13,092 clips) splits. Each clip includes start and end timestamps, a short textual narration, and a verb and noun class that correspond to the narration. There are 3805 action classes, 97 verb classes, and 300 noun classes. We evaluate our pre-trained model on the validation split.
EGTEA egtea. EGTEA comprises 28 hours of egocentric cooking videos, annotated with 10,321 instances of fine-grained actions across 106 classes. The average action duration is 3.2 seconds. For our experiments, we use only the visual frames as input. We follow prior works kazakos2021little; lavila and report top-1 accuracy and mean class accuracy on all three test splits, including 2,022 testing instances for each split.
H2O Kwon2021h2o. H2O is a dataset capturing egocentric hand-object interactions in a laboratory, including 36 action classes. The egocentric data is captured from an Azure Kinect camera mounted egocentrically for recordings. Since our primary target is to evaluate the transfer learning capability of visual representation, the train/val splits have 7862/11638 frames.
A.3 Experimental Details
Zero-Shot Video-Text Retrieval in EK100MIR and EgoMCQ. We perform video-text matching with 16 frames as input for EK100MIR and 4 frames for EgoMCQ, following zhao2023avion.
Zero-Shot Action Recognition in EGTEA. We follow the evaluation protocol proposed by egtea to compute the mean performance across all evaluation splits. This involves performing video-text retrieval between video clips and the action text labels, which are prompted by prepending ”#C C …”. During inference, we apply three spatial crops of size $224\times 224$ from each $256\times 256$ frame of 10 video clip, averaging predictions across these crops to produce the final results.
Depth Estimation in H2O. The frozen visual encoder produces feature maps of dimension 768, which are passed to a linear decoder to estimate depths with a resolution of $720\times 720$. The model is trained for 10 epochs with a batch size of 64 and a learning rate of 0.0005, where the first 1.5 epochs serve as a warm-up phase. Our evaluation code is built upon Probing3D cvpr-2024probing3D.
Robot Manipulation in Franka Kitchen. In Franka Kitchen environment, all baselines apply imitation learning for visuomotor control. A policy network is trained for each task using observations from the environment and video representation from EgoDTM. To adapt EgoDTM and LaViLa, we repeat the image observation 4 times as video input. For each task, the experiment is conducted from two different camera viewpoints for two random seeds using 50 randomly sampled trajectories. The final result is the average of the success rate. Our codebase is built upon MPI zeng2024mpi.
Natural Language Queries on Ego4D. The task typically operates on a 6-minute video. Using the video compression technique from caba2015activitynet, we compress the original video frames by 6 times to save storage. Then we extract the features with the fps of 1.87 and sampling frame number 4. We takes 256 dimension global video features and 768 dimension BERT features as input. Our codebase is built upon EgoVLP lin2022egovlp.
Appendix B Additional Quantitative Results
@@
Appendix C Additional Qualitative Results
LLM Prompts. We use the LLM prompts in Figure 11 to enrich the texts with HOI shape and movement information. Specifically, the shape information is provided by the HOI mask areas, where the large, medium, small object occupies [0.1,1], [0.01, 0.1] and [0, 0.01] areas, respectively.
Generated Data. Examples of generated HOI boxes, HOI masks, depths, and enriched texts are illustrated in Figure 12.
@@ Appendix C Additional Qualitative Results
Appendix D Discussions
Comparison with Related Works that Pretrained with Depth. ImageBind girdhar2023imagebind and LanguageBind zhu2024languagebind are the most relevant depth-based vision-language pretraining works. These methods employ multiple encoders to align various modalities within a unified feature space through contrastive learning. While both methods utilize depth for pretraining, their application may be less impactful when applied to conventional third-person datasets. In contrast, depth is essential for egocentric perception, where spatial awareness is critical for understanding human indoor activities. Furthermore, their pretraining processes treat depth as an input rather than a prediction target, resulting in depth features that lack pixel-level 3D information and video representations that remain unaware of 3D structure. In our work, we aim to enable video representations to predict depth maps, thereby embedding 3D awareness directly into the representations.
Potential for Real-World Applications. While our model demonstrates improvements over text-pretrained models in both video understanding and robotic manipulation tasks, it falls short of state-of-the-art performance of manipulation models. However, our model predicts more meaningful depth maps in real-world settings than in simulations, offering promising potential for real-world deployment.
Figure 11: LLM prompt strategy for generating enriched text from HOI masks and the original text.
Figure 12: Illustration of data generated by our data generation pipelines, including intermediate HOI boxes, masks, and the enriched texts and depth maps used as supervision signals. The text that includes HOI movements is marked blue, while the contents that include HOI spatial information are marked red.